# How does the optimizer implicitly bias the model merging loss landscape?

### v1

@@ 6 Conclusion
We study how optimizer choices implicitly shape the model-merging loss landscape and highlight the effective noise scale as a unifying factor. Learning rate, weight decay, batch size, and data augmentation all modulate this noise, which in turn determines whether independently trained solutions are compatible for merging. The relationship is non-monotonic – too little noise yields incompatible solutions, too much destabilizes training, but an intermediate “sweet spot” enables effective merging. These results extend prior works connecting noise to flatness and generalization of individual models, showing that noise also shapes the compatibility across independent solutions. Future work should explore how to further tune effective noise, architectural choices, and pretraining strategies to train models that are not only strong individually but also inherently mergeable. To summarize our contributions in one sentence: tune the noise to tune merging effectiveness.
Limitations. No new theoretical guarantees are developed, and no truly large-scale experiments are conducted due to our limited computational resources. We studied the standard merging methods, that form the foundation of state-of-the-art approaches. Our goal was to use a set of simple, diverse, but realistic experimental setups to understand the role of optimization in model merging.
@@ Acknowledgements
Appendix A Detailed experiment setting
@@ B.3 Transfer learning: ViT, ConvNext-T
Figure 12: Larger learning rate enables easier merging under transfer learning for both ID and OOD datasets. The pretrained architecture is ConvNext-T trained on IN1k and finetuned on CIFAR10. The evaluation is done on the test set ID of CIFAR10 and test set OOD CIFAR10.1 (Recht et al., 2018).
@@ B.4 Language modeling
B.5 Training loss of decayed models
@@
B.6 Merging fails due to high effective noise
@@ Appendix D Additional results on loss landscape
D.1 Transition phase: hills, flatland, and valleys
@@ D.2 Flatness
D.3 Landscape vs. effective noise
(a) Low noise.
(b) Optimal noise (“sweet spot“).
(c) High noise.
Figure 27: 2D loss slices in the plane spanned by the base model and two fine-tuned models under varying effective noise scales. Low and high noise both yield broad, flat corridors between the fine-tuned solutions, whereas an intermediate (optimal) noise level introduces performance gains (“valleys”) between them. Model: ResNet18; dataset: CIFAR100.
Appendix E LLM research assistance
We use LLM to assist this research project in the following tasks: manuscript polishing and retrieval of related work. For both tasks, we make mild use of LLM for the manuscript writing phase. In particular, polishing has been used only to improve the flow of the sentences, while the retrieval of contents has been used to find a few related works.

### v2

@@ 7 Conclusion
We study how optimizer choices implicitly shape the model merging loss landscape and highlight the effective noise scale as a unifying factor. Learning rate, weight decay, batch size, and data augmentation all modulate this noise, which in turn determines the compatibility of independently trained solutions. This relationship is non-monotonic – too little noise yields low benefits, too much destabilizes training, but an intermediate “critical point” enables effective merging. In practice, mergeability appears to be primarily determined by effective noise levels, suggesting that hyperparameter search can be simplified by focusing on fewer dimensions.
Our findings extend prior work connecting optimization trajectory noise to flatness and generalization of individual models, showing that noise also shapes the compatibility of independent solutions. However, many open questions remain. For example, how can we systematically tune effective noise levels, architectural designs, and pretraining strategies to produce models that are not only strong individually but also inherently mergeable with other solutions?
Limitations. No theoretical guarantees are developed, and no truly large-scale experiments are conducted due to our limited computational resources. We studied only the standard merging methods, which form the foundation of the latest approaches. Our goal was to use a set of simple, diverse, but realistic setups to understand the role of optimization in model merging.
@@ Acknowledgements
Author contributions
Chenxiang Zhang led the project, contributing the majority of experiments, infrastructure, figures, framing, direction, and writing. Alexander Theus conceived the effective noise scale framework, designed and ran the experiments validating it, and contributed to figures, framing, and writing. Antonio Orvieto contributed to the framing, direction, and writing. Damien Teney, Jun Pang, and Sjouke Mauw contributed to the framing and direction.
Appendix Contents
Appendix A Experiment setting details
@@ B.4 Language modeling
B.5 Effective noise for DenseNet101 on Tiny-ImageNet
Figure 14: Effective noise scale controls merging effectiveness across batch sizes $\{16,32,64,128,256\}$ and learning rates $\{10^{-3},10^{-2},3\cdot 10^{-2},5\cdot 10^{-2}\}$ for SGD with momentum $0.9$ and weight decay $5\cdot 10^{-4}$. A shared base model is trained for 200,000 steps; at 61 checkpoints along this trajectory, two independent seeds are fine-tuned for 12,000 further steps with a decayed learning rate and merged via uniform weight averaging, yielding 1,220 merged models in total. DenseNet-121 trained on Tiny-ImageNet.
B.6 Training loss of decayed models
@@
B.7 Merging fails due to high effective noise
@@ C.4 Merging different tasks
Figure 25: TIES merging of two different tasks across $\alpha$ values. Similar setups (antidiagonal) consistently have better merged models.
@@ Appendix D Additional results on loss landscape
D.1 Permutation symmetries
(a) Epoch 200
(b) Epoch 200
(c) Epoch 200
(d) Epoch 800
(e) Epoch 800
(f) Epoch 800
(g) Epoch 1600
(h) Epoch 1600
(i) Epoch 1600
Figure 26: Larger learning rate $\eta$ identifies a broader basin that simplify model permutation across different training phase. The setup uses ResNet18 trained on CIFAR100 as in Section 3. The $x$ and $y$-axis are normalized to enable a visual comparison of the basin size.
Figure 27: Accuracy gain after permutation invariance for CIFAR100. We validate the following hypothesis: the minima identified with a larger effective noise simplifies the rebasing process. We train two sets of models $\bm{\theta}_{A}$ and $\bm{\theta}_{B}$ with random initialization using the same setup as in Section 3.2. The weight-based matching is used to rebase as $\pi(\bm{\theta}_{A})\approx\bm{\theta}_{B}$. Then, we measure the permutation invariant $gain_{inv}=acc(\pi(\bm{\theta}_{A}))-acc(\bm{\theta}_{B})$. A larger $gain_{inv}$ value corresponds to a more advantageous basin for merging. We observe a clear correlation between the standard merging gaingain obtained in the shared initialization and branching setup ($y$-axis) against the merging $gain_{inv}$ between independent initialized models ($x$-axis). In particular, a larger lr (or effective noise) helps to identify “flatter” basins that also enables more effective rebasin.
(a) Vanilla averaging
(b) Weight matching
(c) Learned Matching
Figure 28: Loss-barrier analysis for small GPT-2 models trained on WikiText-103 under the alignment methods of Theus et al. (2025). Panels (a)–(c) show the complete loss interpolation curves for different learning rates, while the remaining panels highlight the peak (maximum) loss barrier extracted from each trajectory. Lower loss barriers are better.
In Figure 28, we evaluate the connectedness of small GPT-2 models trained from scratch on WikiText-103 under varying learning rates. All models are 6-layer GPT-2–style decoders (block size 512, $d_{\text{model}}=512$, $n_{\text{head}}=8$, $n_{\text{inner}}=2048$), trained with the GPT-2 tokenizer using a batch size of 32 for 10 epochs, weight decay $0.01$, and a learning-rate warmup ratio of $0.05$. To obtain optimal neuron alignments, we apply the symmetry-aware merging methods of Theus et al. (2025). We consider three settings: vanilla averaging (no alignment), weight matching (alignment via maximizing parameter similarity), and learned matching (alignment optimized directly for next-token prediction on WikiText-103). As in our experiments without symmetry alignment, larger learning rates tend to improve connectivity and reduce loss barriers. However, consistent with prior observations that text-based Transformers trained from scratch do not exhibit linear mode connectivity, we see no cases where interpolation reduces the loss.
D.2 Transition phase: hills, flatland, and valleys
@@ D.3 Flatness
D.4 Landscape vs. effective noise
(a) Low noise
(b) Optimal noise
(c) High noise
Figure 32: ResNet18 trained on CIFAR100 2D loss slices in the plane spanned by the base model and two fine-tuned models under varying effective noise scales. Low and high noise both yield broad, flat corridors between the fine-tuned solutions, whereas an intermediate (optimal) noise level introduces performance gains between them.
Appendix E Additional results for training dynamics
E.1 Scheduler
(a) CIFAR10
(b) CIFAR100
Figure 33: Comparison between WSD and cosine scheduler.
We use the same setup described in Appendix A, varying only the scheduler for the whole training duration. The same training budget (epochs) is used. Figure 33 shows that WSD scheduler enables easier merging, especially when bifurcating in the final phase where the learning rate of cosine is already small.
E.2 Optimizer
(c) AdamW
(d) SGD
Figure 34: Optimizer comparison in transfer learning. Using AdamW, larger learning rate enables easier merging for both ID and OOD datasets, while for SGD, benefits are only for ID dataset. The pretrained architecture is ConvNext-T trained on IN1k and finetuned on CIFAR10. The test set ID is CIFAR10 and test set OOD is CIFAR10.1 (Recht et al., 2018).
We compare how the optimizer affects the merging effectiveness between AdamW and SGD. Note that in this experiment, SGD with small lr required $20\times$ more steps compared to AdamW for convergence to near zero training loss. Figure 34 shows that SGD-trained models have larger performance gain with larger lr for ID dataset, but not for OOD dataset. Moreover, SGD trained models have a lower final performance compared to AdamW models ($95\%$ vs $96\%$).
Appendix F Additional results on TIES merging
Figure 35: TIES merging of models trained on three different tasks (RESISC45, FMoW, CIFAR10). Results are averaged over three seeds.
We study whether more advanced merging methods can reduce sensitivity to hyperparams as suggested. First, we apply TIES directly to the existing setting in Section 4 with two tasks (RESISC45, FMoW). We use TIES to keep $70\%$ of the values and “mean” aggregation.
TIES merging can partially counteract the high noise. We extend the setting in Section 4.1 by applying TIES to three tasks (RESISC45, FMoW, and CIFAR10) and measure its normalized accuracy across interpolation. Figure 35 shows that at small $\alpha<0.5$, a larger lr trained models have the highest performance, while at a larger $\alpha>0.5$, larger lr becomes unstable. This follows the same trend observed in the main sections.
Appendix G Additional results on SWA
(a) $k=3$
(b) $k=10$
(c) $k=3$
(d) $k=10$
Figure 36: Accuracy gain when applying stochastic weight averaging (SWA).
Using the models trained in Section 3.2, we apply stochastic weight averaging (SWA) to the last $k$ checkpoints of the branched models, obtaining $\theta^{swa}_{A}$ and $\theta^{swa}_{B}$, which are merged into $\theta^{swa}_{m}$. We define $gain_{swa}=acc(\theta^{swa}_{m})-0.5*(acc(\theta^{swa}_{A})+acc(\theta^{swa}_{B}))$ to measure the accuracy gain of SWA models after merging.
Figure 36 (top) shows that SWA endpoints can also benefit from merging the branched models $\theta^{swa}_{A}$ and $\theta^{swa}_{B}$. However, the merge gains are lower compared to the standard setting w/o SWA. This is because SWA already incorporates the benefit of large lr (noise) to explore wider valleys by merging the models along the same trajectory, while merging combines models from different trajectories. Figure 36 (bottom) shows that the final accuracies are comparable, and the methods are complementary. These results support the conclusion that effective noise governs mergeability, including SWA.
Appendix H LLM research assistance
We use LLM to assist this research project in the following tasks: manuscript polishing and retrieval of related work. For both tasks, we make mild use of LLM for the manuscript writing phase. In particular, polishing has been used only to improve the flow of the sentences, while the retrieval of contents has been used to find related works.